# Should We Still Pretrain Encoders with Masked Language Modeling?

### v3

@@ Abstract
Learning high-quality text representations is fundamental to a wide range of NLP tasks. While encoder pretraining has traditionally relied on Masked Language Modeling (MLM), recent evidence suggests that decoder models pretrained with Causal Language Modeling (CLM) can be effectively repurposed as encoders, often surpassing traditional encoders on text representation benchmarks. However, it remains unclear whether these gains reflect an inherent advantage of the CLM approach or arise from confounding factors such as model and data scale. In this paper, we address this question through a series of large-scale, carefully controlled pretraining ablations, training a total of 38 models ranging from 210 million to 1 billion parameters, and conducting over 15,000 fine-tuning and evaluation runs. We find that while training with MLM generally yields better performance across text representation tasks, CLM-trained models are more data-efficient and demonstrate improved fine-tuning stability. Building on these findings, we experimentally show that a biphasic training strategy that sequentially applies CLM and then MLM, achieves optimal performance under a fixed computational training budget. Moreover, we demonstrate that this strategy becomes more appealing when initializing from readily available pretrained CLM models, reducing the computational burden needed to train best-in-class encoder models. We release all project artifacts at https://hf.co/MLMvsCLM to foster further research.
@@ 1 Introduction
In this paper, we present a controlled study examining the impact of training with MLM and CLM on learning textual representations. We compare models with equal architecture and size, trained on equal amount of data, and evaluate them on an extensive range of text representation tasks. We start by comparing pretraining exclusively with either approach, and then investigate a two-stage pretraining alternative, where CLM and MLM are applied sequentially. Additionally, leveraging existing pretrained models, we investigate a continued pretraining setup, where training starts from models initially trained with either MLM or CLM and is then extended with MLM for a portion of the pretraining steps. In total, we train 38 models ranging from 210 million to 1 billion parameters and conduct over 15,000 fine-tuning and evaluation runs to ensure fair design comparisons, amounting to 110k MI250X GPU hours. These large-scale experiments show that:
@@ 2 Experimental Setup
Figure 1: Experimental setup overview and key results on sequence classification (610M model size, 40% MLM ratio). The upper plot shows SC downstream performance as a function of the CLM–to-MLM step ratio during pretraining. The lower plot illustrates the effect of applying MLM CPT to models initially trained with either MLM or CLM (Base).
@@
  • PFS experiments explore various CLM-to-MLM step ratios (0%–100%, 25%–75%, 50%–50%, 75%–25%, and 100%–0%) across 3 training lengths (12,000, 22,000, and 42,000 steps), focusing on the 610M model size with a 40% masking ratio for MLM. This setup is also used to evaluate the effect of masking ratio on models pretrained with CLM followed by MLM. To minimize redundant computation, we reuse checkpoints from the base pretraining runs whenever possible, applying learning rate decay only during the final 2,000 steps. In total, these runs result in 17 new models and account for 120,000 additional training steps and 16k GPU hours.
@@ 3 Pretraining with CLM or MLM
Bidirectional attention during pretraining tends to enhance representation quality on downstream tasks (Figure 2). The performance gap between models trained with MLM and those trained with CLM is especially noticeable on SC and QA tasks, and remains consistent across model sizes. The largest discrepancy is observed on QA, which appears particularly sensitive to the absence of bidirectional attention during pretraining. Interestingly, some task-specific trends emerge: for instance, the MLM–to-CLM gap widens with increasing model size on SC, but narrows on IR.
@@ 7 Conclusion
Our study challenges the long-standing assumption that Masked Language Modeling (MLM) is the universally optimal approach for encoder pretraining. Through large-scale experiments at the 100B training token scale, we provide robust empirical evidence that encoders can be trained more data-efficiently without relying solely on MLM. In particular, we demonstrate that decoder-only models trained with Causal Language Modeling (CLM) offer clear advantages in data efficiency, training stability, and performance on specific text representation tasks. Moreover, we show that sequential training (first with CLM followed by MLM) is more effective than MLM alone. These results suggest that the most effective path to high-performing encoder models may involve leveraging pretrained decoder models and continuing with MLM-based training, paving the way for future state of the art encoders at a discounted price. We hope our findings and released resources will support further advancements in this area.
@@ 8 Limitations and Future Work
Our study deliberately focuses on a targeted set of parameters, namely learning objective, model size, training setup, data budget, and masking ratio. As in large-scale studies such as (Kaplan et al., 2020; Hoffmann et al., 2022), this required fixing several other parameters, including the model architecture, tokenizer, language, and training mixture. While these factors are certainly of interest, varying them would introduce significant computational overhead, making them difficult to incorporate into our current scope. We therefore view them as a promising direction for future work, including extensions to multilingual settings or to additional modalities such as vision–language.
@@ Appendix B Details on Evaluation Datasets
All SC datasets are sourced form the GLUE (Wang et al., 2018) benchmark.
@@ C.3 CLM-to-MLM Ratio
This appendix provides detailed results in the PFS setting, analyzing the impact of varying the CLM–MLM step ratio during pretraining. Results are presented for 610M models with 40% masking for MLM. Table 13, Table 11, Table 12, and Table 13 present dataset-level results for SC, TC, QA, and IR, respectively.
@@ Appendix D Additional Results
Figure 10: Downstream performance with respect to pretraining FLOPs for CLM and MLM. Results are reported for 610M models, with a 40% masking ratio for MLM. M. Training FLOPs are computed following the formula provided by Hoffmann et al. (2022).
Figure 11: Impact of two-stage CLM+MLM pretraining on downstream performance after 42,000 training steps. Results are reported for the 1B model, with MLM applied using a 40% masking ratio. To keep computation costs manageable, CLM+MLM results are shown only for the 25%–75% regime.

### v4

@@ Abstract
Learning high-quality text representations is fundamental to a wide range of NLP tasks. While encoder pretraining has traditionally relied on Masked Language Modeling (MLM), recent evidence suggests that decoder models pretrained with Causal Language Modeling (CLM) can be effectively repurposed as encoders, often surpassing traditional encoders on text representation benchmarks. However, it remains unclear whether these gains reflect an inherent advantage of the CLM approach or arise from confounding factors such as model and data scale. In this paper, we address this question through a series of large-scale, carefully controlled pretraining ablations, training a total of 38 models ranging from 210 million to 1 billion parameters, and conducting over 15,000 fine-tuning and evaluation runs. We find that while training with MLM generally yields better performance across text representation tasks, CLM-trained models are more data-efficient and demonstrate improved fine-tuning stability. Building on these findings, we experimentally show that a biphasic training strategy that sequentially applies CLM and then MLM achieves optimal performance under a fixed computational training budget. Moreover, we demonstrate that this strategy becomes more appealing when initializing from readily available pretrained CLM models, reducing the computational burden needed to train best-in-class encoder models. We release all project artifacts at https://hf.co/MLMvsCLM to foster further research.
@@ 1 Introduction
In this paper, we present a controlled study examining the impact of training with MLM and CLM on learning textual representations. We compare models with equal architecture and size, trained on an equal amount of data, and evaluate them on an extensive range of text representation tasks. We start by comparing pretraining exclusively with either approach, and then investigate a two-stage pretraining alternative, where CLM and MLM are applied sequentially. Additionally, leveraging existing pretrained models, we investigate a continued pretraining setup, where training starts from models initially trained with either MLM or CLM and is then extended with MLM for a portion of the pretraining steps. In total, we train 38 models ranging from 210 million to 1 billion parameters and conduct over 15,000 fine-tuning and evaluation runs to ensure fair design comparisons, amounting to 110k MI250X GPU hours. These large-scale experiments show that:
@@ 2 Experimental Setup
Figure 1: Experimental setup overview and key results on sequence classification (610M model size, 40% MLM ratio). The upper plot shows SC downstream performance as a function of the CLM-to-MLM step ratio during pretraining. The lower plot illustrates the effect of applying MLM CPT to models initially trained with either MLM or CLM (Base).
@@
  • PFS experiments explore various CLM-to-MLM step ratios (0%-100%, 25%-75%, 50%-50%, 75%-25%, and 100%-0%) across 3 training lengths (12,000, 22,000, and 42,000 steps), focusing on the 610M model size with a 40% masking ratio for MLM. This setup is also used to evaluate the effect of masking ratio on models pretrained with CLM followed by MLM. To minimize redundant computation, we reuse checkpoints from the base pretraining runs whenever possible, applying learning rate decay only during the final 2,000 steps. In total, these runs result in 17 new models and account for 120,000 additional training steps and 16k GPU hours.
@@ 3 Pretraining with CLM or MLM
Bidirectional attention during pretraining tends to enhance representation quality on downstream tasks (Figure 2). The performance gap between models trained with MLM and those trained with CLM is especially noticeable on SC and QA tasks, and remains consistent across model sizes. The largest discrepancy is observed on QA, which appears particularly sensitive to the absence of bidirectional attention during pretraining. Interestingly, some task-specific trends emerge: for instance, the MLM-to-CLM gap widens with increasing model size on SC, but narrows on IR.
@@ 7 Conclusion
Our study challenges the long-standing assumption that Masked Language Modeling (MLM) is the universally optimal approach for encoder pretraining. Through large-scale experiments at the 100B training token scale, we provide robust empirical evidence that encoders can be trained more data-efficiently without relying solely on MLM. In particular, we demonstrate that decoder-only models trained with Causal Language Modeling (CLM) offer clear advantages in data efficiency, training stability, and performance on specific text representation tasks. Moreover, we show that sequential training (first with CLM followed by MLM) is more effective than MLM alone. These results suggest that the most effective path to high-performing encoder models may involve leveraging pretrained decoder models and continuing with MLM-based training, paving the way for future state-of-the-art encoders at a discounted price. We hope our findings and released resources will support further advancements in this area.
@@ 8 Limitations and Future Work
Our study deliberately focuses on a targeted set of parameters, namely learning objective, model size, training setup, data budget, and masking ratio. As in large-scale studies such as (Kaplan et al., 2020; Hoffmann et al., 2022), this required fixing several other parameters, including the model architecture, tokenizer, language, and training mixture. While these factors are certainly of interest, varying them would introduce significant computational overhead, making them difficult to incorporate into our current scope. We therefore view them as a promising direction for future work, including extensions to multilingual settings or to additional modalities such as vision-language.
@@ Appendix B Details on Evaluation Datasets
All SC datasets are sourced from the GLUE (Wang et al., 2018) benchmark.
@@ C.3 CLM-to-MLM Ratio
This appendix provides detailed results in the PFS setting, analyzing the impact of varying the CLM-MLM step ratio during pretraining. Results are presented for 610M models with 40% masking for MLM. Table 13, Table 11, Table 12, and Table 13 present dataset-level results for SC, TC, QA, and IR, respectively.
@@ Appendix D Additional Results
Figure 10: Downstream performance with respect to pretraining FLOPs for CLM and MLM. Results are reported for 610M models, with a 40% masking ratio for MLM. Training FLOPs are computed following the formula provided by Hoffmann et al. (2022).
Figure 11: Impact of two-stage CLM+MLM pretraining on downstream performance after 42,000 training steps. Results are reported for the 1B model, with MLM applied using a 40% masking ratio. To keep computation costs manageable, CLM+MLM results are shown only for the 25%-75% regime.